# What Did My Car Say? Autonomous Vehicle Explanation Errors, Context, and Personal Traits Impact Comfort, Reliance, Satisfaction, and Driving Confidence

### v1

@@ Abstract.
ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing User studies
@@ 1. Introduction
However, although potentially increasing trust, explanation of AI interactions in the real world can contain errors. Prior work has shown that when autonomous vehicles exhibit driving errors, passenger trust and willingness to rely on the vehicle can deteriorate quickly (Seet et al., 2020; Kaplan et al., 2023). Far less is known about the impact of explanation errors (Cabitza et al., 2024; Kenny et al., 2021). Particularly for safety-critical systems like autonomous vehicles – that may rely on explanations and other in-vehicle communications to elicit trust comfort, and safe reliance with users – knowing the consequences of errors is pivotal to safe deployment. Understanding these consequences is essential for user reliance because, without this knowledge, AV systems cannot be deployed safely or ethically (Martinho et al., 2021).
@@
In addition, human-AV interactions do not exist in a vacuum: they are sensitive to the contextual demands of the external driving environment (Hoff and Bashir, 2015). It is important to consider the driving context in which explanation errors occur, as prior work shows that context may significantly influence how people interact with AI-based systems (Lim et al., 2009; Schilit et al., 1994), including autonomous vehicles (Capallera et al., 2022; De Salis et al., 2020). The complexity of a driving situation and perceived risk of harm have been singled out as particular factors of interest (Ha et al., 2020; Kaufman et al., 2024b), as these might drive a person’s explanatory needs and reliance behaviors.
@@
Finally, recent work has posited that AV explanations should be tailored to meet the specific needs of the people interacting with the system (Ma and Feng, 2023; Kaufman et al., 2024b). It is well known that people may interact differently with AI-based systems (Schneider and Handali, 2019) based on the individual characteristics or prior experiences they may have (Ayoub et al., 2021b). Particularly well studied are domain expertise (Araujo et al., 2020; Pazzani et al., 2022; Kaufman and Kirsh, 2022) and initial (dispositional) trust (Hoff and Bashir, 2015).
@@
In the present study, we examine the impact of explanation errors across a variety of realistic, simulated driving scenarios. To deepen our investigation and understand the significance of the type of error presented, participants were shown AV explanations at three distinct accuracy levels. We measured the effect of explanation errors on four main outcomes: comfort relying on the AV, preference for control, confidence in the AV’s driving ability, and explanation satisfaction. We include measures of scenario context – perceived harm and driving difficulty – to assess their impact on our driving outcomes, including how they may moderate the relationship between explanation errors and user perceptions. To explore the influence of individual differences, we also included measures of trust and expertise to determine if they predict study outcomes.
@@
Our findings reveal that explanation errors, contextual characteristics, and personal traits significantly impact how a person may think, feel, and behave towards AVs. Explanation errors negatively affected all outcomes, with impacts proportional to the magnitude and negative implications of the error. Harm and driving difficulty directly impacted outcomes as well as moderated the relationship between errors and outcomes, though in opposing ways. Overall, harm was generally seen as more important and more negative than difficulty. Participants with higher expertise tended to trust AVs more, and these each correlated with more positive outcome ratings in turn.
In sum, we contribute:
  • A study on the impact of explanation errors based on information type (‘why’ and ’what + why’).
  • An assessment of how contextual harm and driving difficulty, initial trust, and expertise impact study outcomes.
  • A set of future design and research implications based on the present findings.
@@ 2.1. AV Trust and Explainability
Human-Centered Explainable AI – Continuous development of explainable AI (XAI) has brought major progress in the quest for trust through AI transparency (Miller, 2019). Approaches to XAI vary, often limited by the availability of the model (Simonyan et al., 2013; Ribeiro et al., 2016). Recent work found that – even when an explanation is given – trust and engagement with the system may not improve unless the right information is given in the right way, at the right time (Wang et al., 2019; Liao and Varshney, 2021). Several studies and theory pieces have demonstrated the value of human-interpretable explanations on user understanding and trust  (Holzinger et al., 2019; Soltani et al., 2022). In particular, explanations that are modeled off those given by human experts have suggested as a means to increase understanding for experts and novices alike (Pazzani et al., 2022; Kaufman and Kirsh, 2022). There have likewise been calls for the need for explanations to be sensitive to particular user characteristics (such as personal traits or experiences)  (Ehsan and Riedl, 2020; Ehsan et al., 2021; Kaufman and Kirsh, 2023) or context of use (including the specific use environment and goals of a user) (Schilit et al., 1994; Kaufman et al., 2024b). Recent discussion on XAI usefulness had proposed that explanations should move beyond simply conveying algorithmic mechanisms for decision-making and instead supporting user hypothesis generation (Miller, 2023). Despite the large amount of research on explanation design, little has been done to understand what happens when these explanations fail. The study we present here is a first step towards filling this knowledge gap, using autonomous vehicles as a specific domain of interest.
@@
Explanation Content – Choosing the appropriate content for an AI explanation is crucial for enhancing rider trust and reliance (Miller, 2019). Recent work as focused on providing a description of what a vehicle is doing and why a vehicle is doing it, as these enable a user to create a momentary evaluation of the AV’s behavior in terms of reliance (Hoff and Bashir, 2015). For example, Koo et al. (2015) present ’how’, ‘why’, and a combination of both in various simulated autonomous driving scenarios to assess driver attitudes and safety performance. They found improved safety when both ’how’ and ‘why’ were presented to drivers, but a preference for ‘why’ explanations alone. Kaufman et al. (2024a) leveraged auditory and visual explanations to teach humans to be better drivers via an ’AI coach’, highlighting the additive value of what and why-type information for transferring knowledge, adding that too much information can cause overwhelm. They emphasize the need for explanations to strike a balance between complexity and comprehensiveness. The impact of what (similar to Koo’s how) and why-type explanation errors – explored in the present study – remains unknown.
Factors Impacting Trust – Knowing what factors may impact a person’s trust and reliance decisions is vitally important to designing XAI explanations to support them. Hoff and Bashir’s theoretical model of trust in automation highlights the importance of three distinct yet interdependent facets of trust: dispositional (e.g. personality or cultural attitudes), situational (e.g. based on a particular context of use), and learned trust (e.g. based on a present evaluation of system performance) (Hoff and Bashir, 2015). Additionally, Kaufman et al. (2024b) developed a framework to understand situational awareness in joint action between humans and AV, a key focus of explainable AI transparency in safety-critical situations like driving. They describe how communications like AV explanations can enable human-AV teamwork to achieve particular goals like safe and trustworthy driving. Factors of interest include external driving conditions, human traits and abilities, and communication preferences and goals – all of which are crucial in managing driving difficulty and reducing the risk of harm. Using these system-based models as a backdrop, AI explanation designers can form hypotheses on how to build more trustworthy systems. In the present study, we build off these models by investigating specific driving context factors and personal traits which may impact a person’s reliance judgments and, in the case of context, how much an error matters.
@@ 2.2. AI Errors
Widespread integration of AI systems into everyday tasks has demonstrated huge benefits to productivity and optimization in many domains (Fauzi et al., 2023). However, concerns over AI errors remain a major point of contention, particularly as systems proliferate. Examples of errors with real-world consequences include language models’ propensity to hallucinate (Xu et al., 2024) and algorithmic bias that favors men over women used in the hiring process (Dastin, 2018).
@@
Knowledge Gap – Though several prior studies have investigated the consequences of accurate AV explanations, the impact of explanation errors on reliance behavior and related outcomes remains unexplored. Of particular interest to us are cases when the autonomous vehicle’s driving performs properly, as this allows us to separate the impact of driving performance from the impact of explanation performance. This has yest to be explored, and we address this important knowledge gap.
@@ 3. Method
We conducted an online experiment using realistic, simulated driving scenarios of an autonomous vehicle driving through various environments. These ranged from rural to urban driving and from routine navigational challenges like driving around construction cones to challenging situations like avoiding collisions with erratic drivers. During the experiment, participants viewed videos of the scenarios and, after each video, provided ratings related to trust, reliance, satisfaction, and evaluation of the AV.
To test the impact of explanation errors, participants were shown three versions of each scenario. Those with an “accurate” explanation were correctly told ‘what’ the AV was doing and ‘why’ it was doing it. Errors were introduced via a “low” error condition, where the AV correctly explained ‘what’ it was doing but incorrectly explained ‘why’, and a “high” error condition, where both ‘what’ and ‘why’ explanations were incorrect. In all three versions of each scenario, the AV drove identically, so only the explanations differed. The AV drove accurately and lawfully in all cases. Participants were randomly shown videos from 9 of 27 possible scenarios, providing ratings after each video, making 27 total video ratings per participant (9 scenarios X 3 accuracy levels). Scenarios were presented in random order. The 27 possible scenarios are described in Table 1.
Table 1. Scenario Descriptions
Scenario Description
10s1 Left turn in busy intersection, navigating around object in road.
10s2 Slowing to avoid collision between two vehicles ahead of the ego vehicle.
10s4 Parallel parking between two cars on side of busy road.
10s5 Slowing to avoid being sideswiped during turn by a vehicle that crosses into ego vehicle’s lane.
10s6 Sudden stop mid-intersection by large lead vehicle in adverse weather.
10s7 Left turn quickly followed by right merge for quick right turn in urban environment in adverse weather.
10s8 Ego vehicle slows for pedestrian crossing road in non-crosswalk area.
4s1 Reversing in parking lot.
4s2 Overtaking cyclist in suburban environment.
4s3 Ego vehicle must stop quickly for a pedestrian who jumps into the road.
4s4 Merging from far right lane to far left lane on highway to avoid emergency vehicles / car accident.
4s5 Ego vehicle must avoid object that falls off of lead vehicle on highway at night.
4s6 Ego vehicle hydroplanes on wet road on highway, and needs to maintain control.
4s7 Ego vehicle needs to pull over for flat tire at high speed.
4s8 Ego vehicle merges left to enter a busy highway.
5s2 Ego vehicle makes blind turn in intersection due to obstructed view. [removed]
5s3 Ego vehicle turning left and avoids a collision with another vehicle who ran a red light (T-bone).
5s4 Car in parallel lane merges into ego vehicle’s path.
5s5 Ego vehicle needs to navigate around a stopped cyclist. [removed]
5s6 Ego vehicle navigates construction zone at night.
7s1 Hidden stop sign at night.
7s3 Lead vehicle quickly decelerates (brake check)
7s4 Ego vehicle crosses the midline to overtake a slow lead vehicle.
7s5 Ego vehicle needs to slow quickly from high speed for a stopped cyclist around a turn. [removed]
7s6 Vehicle failure (flat tire) in small parking lot.
xs2 Ego vehicle waits for child to cross crosswalk before turning right at stop sign.
xs3 Ego vehicle stops during right turn to avoid collision with vehicle turning right from incorrect (left) lane.
@@
Figure 1. Example images from four driving scenarios: (1) the AV slows for a pedestrian crossing the road on a foggy day in the city, (2) the AV is cut off by a vehicle turning right from the left lane, (3) the AV merges around a slow lead vehicle in a forested area, (4) the AV moves past a cyclist on a suburban street. In the bottom right corner of each scenario, written AV explanations appear on the vehicle’s dashboard display.
[…]
Example images from four driving scenarios: (1) the AV slows for a pedestrian crossing the road on a foggy day in the city, (2) the AV is cut off by a vehicle turning right from the left lane, (3) the AV merges around a slow lead vehicle in a forested area, (4) the AV moves past a cyclist on a suburban street. In the bottom right corner of each scenario, written AV explanations appear on the vehicle’s dashboard display.
@@ 3.2. Simulated Driving Scenarios
All simulated driving videos were custom-made by the research team using DReye VR (Silvera et al., 2022), a tool for creating realistic driving scenarios using the open-source driving simulator CARLA (Dosovitskiy et al., 2017). Scenario design was inspired by the National Highway Traffic Safety Administration (NHTSA) list of common driving situations that result in vehicle crashes (Najm et al., 2007). Examples include unexpected pedestrians in the road, collision avoidance from erratic other drivers, construction and emergency vehicle zones, parallel parking and reversing from park, navigating around stopped or slowed vehicles or cyclists, and dealing with flat tires or hydroplaning. Videos lasted between 10 and 30 seconds each (Examples in Fig. 1).
Data from a total of 27 scenarios (see Table 1) was collected during the study; three scenarios were removed during analysis due to the AV’s driving determined to be imperfect (going above the speed limit, crossing over the center line during a turn, and improper yielding), making a total of 24 included in the final analysis. Data was cleaned prior to analysis to ensure that all videos were viewed in full.

### v3

# What Did My Car Say? Impact of Autonomous Vehicle Explanation Errors and Driving Context On Comfort, Reliance, Satisfaction, and Driving Confidence

@@ Abstract.
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713088isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing User studies
@@ 1. Introduction
However, although potentially increasing trust, explanation of AI behavior and decision-making in the real world can contain errors. Prior work has shown that when autonomous vehicles exhibit driving errors, passenger trust and willingness to rely on the vehicle can deteriorate quickly (Seet et al., 2020; Kaplan et al., 2023). Far less is known about the impact of explanation errors, such as when an AV miscommunicates what it is doing or why (Cabitza et al., 2024; Kenny et al., 2021). Particularly for safety-critical systems like autonomous vehicles – that may rely on explanations and other in-vehicle communications to elicit trust comfort, and safe reliance with users – knowing the consequences of errors is pivotal to safe deployment. We hypothesize that explanation errors, like driving errors, will negatively impact perceptions of an AV. Understanding these consequences is essential for user reliance because, without this knowledge, AV systems cannot be deployed safely or ethically (Martinho et al., 2021).
Autonomous vehicles offer a unique context to study explainable AI errors, as vehicle errors may be easier to detect for lay users with basic driving knowledge compared to errors in more specialized domains, such as medical diagnosis (Rajpurkar et al., 2017; Kaufman and Kirsh, 2023) or bird identification (Pazzani et al., 2022; Soltani et al., 2022), where domain expertise is required for error detection. This distinct characteristic allows non-experts to evaluate AV explanations with relatively high confidence, providing a unique perspective on how explanation errors might influence user perceptions, judgments, and trust in AI. Prior research in XAI suggests that task complexity (Salimzadeh et al., 2023) and domain expertise (Nourani et al., 2020) are crucial factors in how users interpret and respond to AI explanations – we extend this understanding by examining explanation errors in a high-risk, non-expert domain, contributing new insights into how explanation accuracy affects user trust and reliance.
@@
Human-AV interactions do not exist in a vacuum: they are sensitive to the contextual demands of the external driving environment (Hoff and Bashir, 2015). It is important to consider the driving context in which explanation errors occur, as this may significantly influence how people interact with AI-based systems (Lim et al., 2009; Schilit et al., 1994), including AVs (Capallera et al., 2022; De Salis et al., 2020). The complexity of a driving situation and perceived risk of harm have been singled out as particular factors of interest (Ha et al., 2020; Kaufman et al., 2024b), as these might drive a person’s explanatory needs and reliance behaviors. Though presently underexplored, we hypothesize that contextual factors may impact rider perceptions of an AV, particularly when errors are introduced.
@@
Finally, recent work has posited that AV explanations should be tailored to meet the specific needs of the people interacting with the system (Ma and Feng, 2023; Kaufman et al., 2024b). It is well known that people may interact differently with AI-based systems (Schneider and Handali, 2019) based on the individual characteristics or prior experiences they may have (Ayoub et al., 2021b; Kaufman et al., 2024c). Particularly well studied are domain expertise (Araujo et al., 2020; Pazzani et al., 2022; Kaufman and Kirsh, 2022) and initial (dispositional) trust (Hoff and Bashir, 2015); we hypothesize that these will also impact AV perceptions.
@@
In the present study, we examine the impact of explanation errors across a variety of realistic, simulated driving scenarios. To deepen our investigation and understand the significance of the type of error presented, participants were shown AV explanations at three distinct accuracy levels: (1) accurate explanation of the AV’s behavior, (2) accurate explanation of what the AV is doing but incorrect rationale for why, and (3) incorrect explanation of what and rationale of why. We measured the effect of these errors on four main outcomes: comfort relying on the AV, preference for control, confidence in the AV’s driving ability, and explanation satisfaction. We include measures of scenario context – perceived harm and driving difficulty – to assess their impact on our driving outcomes, including how they may moderate the relationship between explanation errors and user perceptions. To explore the influence of individual differences, we also included measures of trust and AV domain expertise to determine if they predict study outcomes.
@@
Our findings reveal that explanation errors, contextual characteristics, and personal traits significantly impact how a person may think, feel, and behave towards AVs. Explanation errors negatively affected all outcomes, with impacts proportional to the magnitude and negative implications of the error. Harm and driving difficulty directly impacted outcomes as well as moderated the relationship between errors and outcomes, though in opposing ways. Overall, harm was generally seen as more important and more negative than difficulty. Participants with higher AV domain expertise tended to trust AVs more, and these each correlated with more positive outcome ratings in turn.
In sum, our research questions are as follows:
  • RQ1: What is the impact of AV explanation errors on participant ratings of an AV, and does the impact differ for errors of different information types (‘why’ errors vs. ‘what and why’ errors)?
  • RQ2: What is the impact of context (perceived harm and driving difficulty) on participant ratings of an AV, and do these contextual factors modify the effect of an error?
  • RQ3: How does participant initial trust and domain expertise impact their ratings of an AV?
  • RQ4: How can findings inform future design and research of explainable AI and autonomous vehicles?
@@ 2.1. AV Trust and Explainability
Human-Centered Explainable AI – Continuous development of explainable AI (XAI) has brought major progress in the quest for trust through AI transparency (Miller, 2019). Approaches to XAI vary, often limited by the availability of the model (Simonyan et al., 2013; Ribeiro et al., 2016). Recent work found that – even when an explanation is given – trust and engagement with the system may not improve unless the right information is given in the right way, at the right time (Wang et al., 2019; Liao and Varshney, 2021). Several studies and theory pieces have demonstrated the value of human-interpretable explanations on user understanding and trust  (Holzinger et al., 2019; Soltani et al., 2022). In particular, explanations that are modeled off those given by human experts have suggested as a means to increase understanding for experts and novices alike (Pazzani et al., 2022; Kaufman and Kirsh, 2022). There have likewise been calls for the need for explanations to be sensitive to particular user characteristics (such as personal traits or experiences)  (Ehsan and Riedl, 2020; Ehsan et al., 2021; Kaufman and Kirsh, 2023) or context of use (including the specific use environment and goals of a user) (Schilit et al., 1994; Kaufman et al., 2024b). Despite the large amount of research on explanation design, little has been done to understand what happens when these explanations fail. The study we present here is a first step towards filling this knowledge gap, using autonomous vehicles as a specific domain of interest.
@@
Explanation Content – Choosing the appropriate content for an AI explanation is crucial for enhancing rider trust and reliance (Miller, 2019). Recent work as focused on providing a description of what a vehicle is doing and why a vehicle is doing it, as these enable a user to create a momentary evaluation of the AV’s behavior in terms of reliance (Hoff and Bashir, 2015). For example, Koo et al. (2015) present ‘how’, ‘why’, and a combination of both in various simulated autonomous driving scenarios to assess driver attitudes and safety performance. They found improved safety when both ’how’ and ‘why’ were presented to drivers, but a preference for ‘why’ explanations alone. Kaufman et al. (2024a) leveraged auditory and visual explanations to teach humans to be better drivers via an ’AI coach’, highlighting the additive value of what and why-type information for transferring knowledge, adding that too much information can cause overwhelm. They emphasize the need for explanations to strike a balance between complexity and comprehensiveness. The impact of what (similar to Koo’s how) and why-type explanation errors – explored in the present study – remains unknown.
Factors Impacting Trust – Knowing what factors may impact a person’s trust and reliance decisions is vitally important to designing XAI explanations to support them. Hoff and Bashir’s theoretical model of trust in automation highlights the importance of three distinct yet interdependent facets of trust: dispositional (e.g. personality or cultural attitudes), situational (e.g. based on a particular context of use), and learned trust (e.g. based on a present evaluation of system performance) (Hoff and Bashir, 2015). Additionally, Kaufman et al. (2024b) developed a framework to understand situational awareness in joint action between humans and AV, a key focus of explainable AI transparency in safety-critical situations like driving. They describe how communications like AV explanations can enable human-AV teamwork to achieve particular goals like safe and trustworthy driving. Factors of interest include external driving conditions, human traits and abilities, and communication preferences and goals – all of which are crucial in managing driving difficulty and reducing the risk of harm. Using these system-based models as a backdrop, AI explanation designers can form hypotheses on how to build more trustworthy systems. In the present study, we build off these models by investigating specific driving context factors (perceived harm and driving difficulty) and personal traits (domain expertise and prior trust) which may impact a person’s reliance judgments and, in the case of context, how much an error matters.
@@ 2.2. AI Errors
Widespread integration of AI systems into everyday tasks has demonstrated huge benefits to productivity and optimization in many domains (Fauzi et al., 2023). However, concerns over AI errors remain a major point of contention, particularly as systems proliferate. Examples of errors with real-world consequences include language models’ propensity to hallucinate (Xu et al., 2024) and algorithmic bias that favors men over women used in the hiring process (Dastin, 2018). In contrast to fields requiring domain-specific knowledge for error detection (e.g., medical diagnosis (Rajpurkar et al., 2017; Kaufman and Kirsh, 2023)), AV explanation errors are often intuitively detectable by lay users. This accessibility may amplify the impact of errors on trust and reliance, as errors are not only observed but judged against personal driving ability.
@@
Knowledge Gap – Though several prior studies have investigated the consequences of accurate AV explanations, the impact of explanation errors on reliance behavior and related outcomes remains unexplored. Of particular interest are cases when the autonomous vehicle’s driving performs properly, as this allows us to separate the impact of driving performance from the impact of explanation performance. We address this knowledge gap.
Figure 1. Example images from four driving scenarios: (1) the AV slows for a pedestrian crossing the road on a foggy day in the city, (2) the AV is cut off by a vehicle turning right from the left lane, (3) the AV merges around a slow lead vehicle in a forested area, (4) the AV moves past a cyclist on a suburban street. Written AV explanations appear on the vehicle’s dashboard display.
[…]
Example images from four driving scenarios: (1) the AV slows for a pedestrian crossing the road on a foggy day in the city, (2) the AV is cut off by a vehicle turning right from the left lane, (3) the AV merges around a slow lead vehicle in a forested area, (4) the AV moves past a cyclist on a suburban street. Written AV explanations appear on the vehicle’s dashboard display.
@@ 3. Method
We conducted an online experiment using realistic, simulated driving scenarios of an AV driving through various environments. These ranged from rural to urban driving and from routine navigational challenges like driving around construction cones to challenging situations like avoiding collisions with erratic drivers. Participants viewed videos of the scenarios and, after each, provided ratings related to trust, reliance, satisfaction, and evaluation of the AV.
To test the impact of explanation errors, participants were shown three versions of each scenario, each differing by the accuracy of the information. This within-subjects design helps us control for individual differences and attribute findings to the error conditions and scenarios themselves. In all three versions of each scenario, the AV drove identically; only the explanations differed. The AV always drove accurately and lawfully. Participants were randomly presented 9 of 27 possible scenarios, resulting in a unique set for each. This approach helped evenly distribute any order effects across participants, reducing systematic bias that could result from a block design, increasing generalizability while preserving a naturalistic flow. Each participant provided 27 total video ratings (9 scenarios X 3 accuracy levels). Scenarios were presented in random order; all had approximately the same number of total ratings. Scenarios are described in Table 1.
Table 1. Scenario Reference Codes and Descriptions
[…]
A list of all of the Scenarios used in the study, including brief descriptions of each. Code Description 10s1 Left turn in busy intersection, navigating around object in road. 10s2 Slowing to avoid collision between two vehicles ahead of the ego vehicle. 10s4 Parallel parking between two cars on side of busy road. 10s5 Slowing to avoid being sideswiped during turn by a vehicle that crosses into ego vehicle’s lane. 10s6 Sudden stop mid-intersection by large lead vehicle in adverse weather. 10s7 Left turn quickly followed by right merge for quick right turn in urban environment in adverse weather. 10s8 Ego vehicle slows for pedestrian crossing road in non-crosswalk area. 4s1 Reversing in parking lot. 4s2 Overtaking cyclist in suburban environment. 4s3 Ego vehicle must stop quickly for a pedestrian who jumps into the road. 4s4 Merging from far right lane to far left lane on highway to avoid emergency vehicles / car accident. 4s5 Ego vehicle must avoid object that falls off of lead vehicle on highway at night. 4s6 Ego vehicle hydroplanes on wet road on highway, and needs to maintain control. 4s7 Ego vehicle needs to pull over for flat tire at high speed. 4s8 Ego vehicle merges left to enter a busy highway. 5s2 Ego vehicle makes blind turn in intersection due to obstructed view. [removed] 5s3 Ego vehicle turning left and avoids a collision with another vehicle who ran a red light (T-bone). 5s4 Car in parallel lane merges into ego vehicle’s path. 5s5 Ego vehicle needs to navigate around a stopped cyclist. [removed] 5s6 Ego vehicle navigates construction zone at night. 7s1 Hidden stop sign at night. 7s3 Lead vehicle quickly decelerates (brake check) 7s4 Ego vehicle crosses the midline to overtake a slow lead vehicle. 7s5 Ego vehicle needs to slow quickly from high speed for a stopped cyclist around a turn. [removed] 7s6 Vehicle failure (flat tire) in small parking lot. xs2 Ego vehicle waits for child to cross crosswalk before turning right at stop sign. xs3 Ego vehicle stops during right turn to avoid collision with vehicle turning right from incorrect (left) lane.
@@ 3.2. Simulated Driving Scenarios
All simulated driving videos were custom-made by the research team using DReye VR (Silvera et al., 2022), a tool for creating realistic driving scenarios using the open-source driving simulator CARLA (Dosovitskiy et al., 2017). Scenario design was inspired by the National Highway Traffic Safety Administration (NHTSA) list of common driving situations that result in vehicle crashes (Najm et al., 2007). Examples include unexpected pedestrians in the road, collision avoidance from erratic other drivers, construction and emergency vehicle zones, parallel parking and reversing from park, navigating around stopped or slowed vehicles or cyclists, and dealing with flat tires or hydroplaning. Videos lasted between 10 and 30 seconds each (Examples in Fig. 1). Three scenarios were removed during analysis due to the AV’s driving determined to be imperfect (going above the speed limit, crossing over the center line during a turn, and improper yielding), making a total of 24 included in the final analysis. Data was cleaned prior to analysis to ensure that all videos were viewed in full.